# CDsampling: An R Package for Constrained D-Optimal Sampling in Paid Research Studies

### v3

@@ 2.1   Constrained lift-one algorithm
In the context of paid research studies, budgetary limitations often necessitate the selection of a subset of participants. We consider the sampling of $n$ subjects from a larger population of $N$, where $n<N$. A typical constraint in such studies is $n_{i}\leq N_{i}$, where $N_{i}$ represents the number of available subjects from the $i$th experimental setting group. This effectively places an upper bound on the sample size for each subgroup (or stratum) ensuring the sampling process does not overdraw from any subgroup. Additional constraints may include $n_{1}+n_{2}+n_{3}+n_{4}\leq 392$ (see the MLM example in Section 3.2), $4n_{1}\geq n_{3}$ (see Example S2.2 in Section S2), etc. The constrained lift-one algorithm seeks the approximate allocation ${\mathbf{w}}$ that maximizes $|{\mathbf{F}}({\mathbf{w}})|$, on a collection of feasible approximate allocations $S\subseteq S_{0}$ where
@@
The set $S$ is presumed to be either a closed convex set or a finite union of closed convex sets. The framework of the constrained lift-one algorithm is provided in Figure 2. The details of the algorithm can be found in Algorithm 1 of Huang et al. (2023). In the constrained lift-one algorithm, the search for the optimal lift-one weight $z$ in step $3^{\circ}$ of Figure 2 is confined within the interval of $[r_{i1},r_{i2}]$ (see Example S2.2 in Supplementary Material and Section S.3 in Huang et al. (2023) for details of finding $[r_{i1},r_{i2}]$). To ensure that the resulting allocation is D-optimal, two additional decision steps, labeled steps $7^{\circ}$ and $8^{\circ}$, are incorporated into the algorithm. The reported allocation in step $10^{\circ}$ is the constrained D-optimal approximate allocation for the study. To illustrate the difference between the lift-one algorithm and the constrained lift-one algorithm, we provide examples in Supplementary Material Section S2.
@@ 2.2   Fisher information for generalized linear models
where $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{m})^{\top}$ is the $m\times p$ design matrix with $\mathbf{X}_{i}=\mathbf{h}(\mathbf{x}_{i})$, $,\mathbf{W}=\text{diag}\{w_{1}\nu_{1},$$\dots,w_{m}\nu_{m}\}$ is a diagonal matrix with $\nu_{i}=\frac{1}{\text{Var}(Y_{i})}(\frac{\partial\mu_{i}}{\partial\eta_{i}})^{2}$. Here $\nu_{i}$ represents how much information the design point $\mathbf{x}_{i}$ contains. The explicit formats of $\nu_{i}$ with different response distributions and link functions can be found in Table 5 of the Supplementary Material of Huang et al. (2023). To calculate the Fisher information matrix $\mathbf{F}$ given the approximate allocation $\mathbf{w}$, we may use the F_func_GLM() function in the CDsampling package. Additionally, the W_func_GLM() function can be used to find the diagonal elements of the matrix $\mathbf{W}$ in the Fisher information matrix (1) of GLM. An example of finding the Fisher information matrix for GLM is provided in Example S1.1 of Supplementary Material.
@@ 2.3   Fisher information for multinomial logistic model
We assume that the responses $\mathbf{Y}_{i}=(Y_{i1},\dots,Y_{iJ})$ follow a multinomial distribution with probabilities $(\pi_{i1},\dots,\pi_{iJ})$, where $\pi_{ij}=P(Y_{i}=j\mid{\mathbf{x}}_{i})$, $Y_{i}\in\{1,\ldots,J\}$ is the $i$th original categorical response, $j=1,\dots,J$, and $\sum_{j}\pi_{ij}=1$. If the response variables are nominal, in other words, there is no natural ordering among different levels, a commonly used MLM is the baseline-category logit model with the $J$th level selected as the baseline category. If the response variable is ordinal, that is, we have a natural ordering of response levels, there are three commonly used MLMs in the literature: cumulative logit model, adjacent logit model, and continuation-ratio logit model (Bu et al., 2020; Dousti Mousavi et al., 2023; Wang and Yang, 2024).
In addition to different logit models, the proportional odds (po) assumption is an important concept in MLMs. The po assumption is a parsimonious model assumption, where a model simultaneously uses all $J-1$ logits in a single model with the same coefficients. This means the covariate effect is constant on the odds ratio among different response levels. When the assumption doesn’t hold, the model is referred to as a non-proportional odds (npo) model and has more parameters in it. When the assumption only holds for part of the parameters, the model is referred to as a partial proportional odds (ppo) model. Commonly used multinomial logit models with po, npo, or ppo assumptions can be summarized in a unified matrix form (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020):
@@
$\mathbf{C}^{\top}=\begin{pmatrix}\mathbf{I}_{J-1}&-\mathbf{I}_{J-1}&\mathbf{0}^{\top}_{J-1}\\ \mathbf{0}^{\top}_{J-1}&\mathbf{0}^{\top}_{J-1}&1\\ \end{pmatrix}$
@@ 3   Examples
Example Specific Finding $I$ set for exact allocation conversion in trial_data and trauma_data examples (see Section 3 and Section S4). iset_func_trauma(), iset_func_trial()
@@ 3.1   Applications in paid research study: $\tt trial\_data$ example
The trial_data dataset is a simulated data for a toy example of paid research studies. This study includes a cohort of $500$ patients for a clinical trial, with gender and age as stratification factors. A logistic regression model incorporates these factors as covariates: gender (coded as 00 for female and $1$ for male) and age (coded as two dummy variables $age\_1$ and $age\_2$ with $(age\_1,age\_2)=(0,0)$ for age group $18\sim 25$; $(1,0)$ for age group $26\sim 64$, and $(0,1)$ for age group $65$ and above). For simplicity, the study assumes binary gender options and a tripartite age categorization. In total, there are $m=6$ combinations of covariate factors. In practice, non-binary gender options and a “prefer not to answer” choice may be included to respect gender diversity and protect patient confidentiality. The response $Y$ denotes the treatment’s efficacy (00 indicating ineffectiveness, $1$ indicating effectiveness). The data is generated by the logistic regression model
@@
Figure 4: The number of patients from different gender (F, M) and age groups ($18-25$, $26-64$, and $\geq 65$) and their responses (00 indicating ineffectiveness, $1$ indicating effectiveness) to treatment in trial_data of CDsampling package.
@@
There are $m=6$ design points, corresponding to gender and age group combinations $(x_{gender\_i},$ $x_{age\_1i},x_{age\_2i})=(0,0,0)$, $(0,1,0)$, $(0,0,1)$, $(1,0,0)$, $(1,1,0)$, and $(1,0,1)$, respectively. The numbers of available volunteers in the six categories are $(N_{1},N_{2},\ldots,N_{6})=(50,40,10,$ 200,150,50)200,150,50)200 , 150 , 50 ). Our goal is to find the constrained D-optimal allocation $(w_{1},w_{2},\dots,w_{6})$ in the set of all feasible allocations $S=\{(w_{1},\ldots,$ $w_{m})^{T}\in S_{0}\mid nw_{i}\leq N_{i},i=1,\ldots,m\}$.
We consider the logistic regression model (3), to find the locally D-optimal design, we assume, for illustrative purposes, that the model parameters $(\beta_{0},\beta_{1},\beta_{21},\beta_{22})$ $=(0,3,3,3)$. We may define the parameters and the model matrix as follows. Subsequently, we find the $\mathbf{W}$ matrix in (1) for the Fisher information matrix $\mathbf{F}$.
@@
  • $\mathbf{w}_{0}$: reports the random initial weights used to initialize the optimization.
@@
In this example, the D-optimal exact allocation is to sample $50$ subjects from the “female, $18-25$” subgroup, $40$ subjects from the “female, $26-64$” subgroup, $10$ subjects from the “female, $\geq 65$” subgroup, and $100$ subjects from the “male, $18-25$” subgroup. Such a design may not explore all the design space, and lead to extreme design cases. In practice, allocating some subjects to the omitted subgroups “male, $26-64$” and “male, $\geq 65$” subgroups could improve the robustness and reduce the risk of overfitting.
Alternatively, one may aim for EW D-optimal allocations when partial coefficients information available with the ${\mathbf{W}}$ matrix substituted by the expectation (EW stands for the expectation of the ${\mathbf{W}}$ matrix). To calculate these expectations, one may define prior distributions for the parameters based on available information. For instance, in this scenario, we assume the following independent prior distributions: $\beta_{0}\sim$ uniform(2,2)22(-2,2)( - 2 , 2 ), $\beta_{1}\sim$ uniform$(-1,5)$, $\beta_{21}\sim$ uniform$(-1,5)$, and $\beta_{22}\sim$ uniform(-1,5). Subsequently, the diagonal elements of ${\mathbf{W}}$ are determined through integration. For $i=1,\dots,m$ and $\eta_{i}=\beta_{0}+x_{gender\_i}\beta_{1}+x_{age\_i1}\beta_{21}+x_{age\_i2}\beta_{22}$, we calculate the key component $E(\nu_{i})$ of the $i$th diagonal element of ${\mathbf{W}}$ through:
[…]
@@
> bounded_uniform(Ni=c(50, 40, 10, 200, 150, 50), nsample=200)Optimal Sampling Results:================================================================================Optimal exact allocation:           F, 18-25 F, 26-64 F, >=65 M, 18-25 M, 26-64 M, >=65allocation 38.0     38.0     10.0    38.0     38.0     38.0--------------------------------------------------------------------------------
@@
To perform a comparison analysis on different sampling strategies, including the constrained D-optimal allocation, the constrained EW D-optimal allocation with uniform priors, the constrained uniform allocation, simple random sample without replacement (SRSWOR), as well as the full data (all the $500$ patients enrolled), we simulate their responses $Y_{ij}$’s based on model (3) with parameter values $(0,3,3,3)$. We apply each sampling strategy to obtain a sample of $n=200$ observations out of $500$, and estimate the parameters using the $200$ observations. The exception is the full data method, where estimation is performed using all $500$ patients. We use the root mean square error (RMSE) to measure the accuracy of the estimates (see Section 4 of Huang et al. (2023) for more theoretical and technical details). We repeat the procedure $100$ times and display the corresponding RMSEs in Figure 5 and simulation codes in Supplementary Material Section S3. Obviously, if we use the full data (all $500$ patients) to fit the model, its RMSE attains the lowest. Besides that, the constrained locally D-optimal allocation and the constrained EW D-optimal allocation have a little higher RMSEs than the full data estimates, but outperform SRSWOR and the constrained uniform allocation. The sampling strategy based on the constrained uniform allocation doesn’t need any model information and is a more robust sampling scheme, which is still better than SRSWOR.
@@ 3.2   Applications in paid research study: $\tt trauma\_data$ example
In this example, we have $m=8$ subgroups, which are combinations of the two covariates categories: dose levels and severity grades. We aim to enroll $n=600$ patients from the $802$ available patients. The collection of feasible allocations is inherently constrained by the number of patients in different severity grades, defined as $S={(w_{1},\ldots,w_{8})^{\top}\in S_{0}\mid n(w_{1}+w_{2}+w_{3}+w_{4})\leq 392,n(w_{5}+w_{6}+w_{7}+w_{8})\leq 410}$. The constraints specify that in the sample, the number of patients with mild symptoms must not exceed $392$ across all dose levels, while those with moderate/severe symptoms must not exceed $410$.
@@
> J=5; p=12; m=8; #response levels; parameters; subgroups> #coefficients> beta = c(-4.047, -0.131, 4.214, -2.225, -0.376, 3.519, -0.302, -0.237,  2.420, 1.386,+ -0.120,  1.284)> #define design matrix of 8 subgroups> Xi=rep(0,J*p*m); dim(Xi)=c(J,p,m);> Xi[,,1] = rbind(c( 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 1, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 1, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,2] = rbind(c( 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 2, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 2, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,3] = rbind(c( 1, 3, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 3, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 3, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,4] = rbind(c( 1, 4, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 4, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 4, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,5] = rbind(c( 1, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 1, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 1, 1, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,6] = rbind(c( 1, 2, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 2, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 2, 1, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,7] = rbind(c( 1, 3, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 3, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 3, 1, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,8] = rbind(c( 1, 4, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 4, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 4, 1, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))
@@
As the trauma_data example doesn’t have bounded constraints, to find the constrained uniform sampling allocation, we use approxtoexact_constrained_func() with one subject in each stratum or subgroup as the input, that is, the approximate allocation $\mathbf{w}=(1/600,1/600,1/600,1/600,$ 1/600,1/600,1/600,1/600)1/600,1/600,1/600,1/600)1 / 600 , 1 / 600 , 1 / 600 , 1 / 600 ). The corresponding $I$ set function is provided in the CDsampling package, and it can be easily defined according to other constraints, see Section S4 of Supplementary Material. Note that the determinant provided by approxtoexact_constrained_func() for different designs are not comparable, as the criteria Fdet_func differ.

### v4

@@ 2.1  Constrained lift-one algorithm
In the context of paid research studies, budgetary limitations often necessitate the selection of a subset of participants. We consider the sampling of $n$ subjects from a larger population of $N$, where $n<N$. A typical constraint in such studies is $n_{i}\leq N_{i}$, where $N_{i}$ represents the number of available subjects from the $i$th experimental setting group. This effectively places an upper bound on the sample size for each subgroup (or stratum) ensuring the sampling process does not overdraw from any subgroup. Additional constraints may include $n_{1}+n_{2}+n_{3}+n_{4}\leq 392$ (see the MLM example in Section 3.2), $4n_{1}\geq n_{3}$ (see Example S2.2 in Supplementary Material Section S2), etc. The constrained lift-one algorithm seeks the approximate allocation $\mathbf{w}$} that maximizes $|{\mathbf{F}}({\mathbf{w}})|$, on a collection of feasible approximate allocations $S\subseteq S_{0}$ where
@@
The set $S$ is presumed to be either a closed convex set or a finite union of closed convex sets. The framework of the constrained lift-one algorithm is provided in Figure 2. The details of the algorithm can be found in Algorithm 1 of Huang et al. (2025). In the constrained lift-one algorithm, the search for the optimal lift-one weight $z$ in step $3^{\circ}$ of Figure 2 is confined within the interval of $r_{i1},r_{i2}]$ (see Example S2.2 in Supplementary Material and Section S.3 in Huang et al. (2025) for details of finding $r_{i1},r_{i2}]$). To ensure that the resulting allocation is D-optimal, two additional decision steps, labeled steps $7^{\circ}$ and $8^{\circ}$, are incorporated into the algorithm. The reported allocation in step $10^{\circ}$ is the constrained D-optimal approximate allocation for the study. To illustrate the difference between the lift-one algorithm and the constrained lift-one algorithm, we provide examples in Supplementary Material Section S2.
@@ 2.2  Fisher information for generalized linear models
where $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{m})^{\top}$ is the $m\times p$ design matrix with $\mathbf{X}_{i}=\mathbf{h}(\mathbf{x}_{i})$, $\mathbf{W}=\text{diag}\{w_{1}\nu_{1},$ $\dots,w_{m}\nu_{m}\}$ is a diagonal matrix with $\nu_{i}=\frac{1}{\text{Var}(Y_{i})}(\frac{\partial\mu_{i}}{\partial\eta_{i}})^{2}$. Here $\nu_{i}$ represents how much information the design point $\mathbf{x}_{i}$ contains. The explicit formats of $\nu_{i}$ with different response distributions and link functions can be found in Table 5 of the Supplementary Material of Huang et al. (2025). To calculate the Fisher information matrix $\mathbf{F}$ given the approximate allocation $\mathbf{w}$, we may use the F_func_GLM() function in the CDsampling package. Additionally, the W_func_GLM() function can be used to find the diagonal elements of the matrix $\mathbf{W}$ in the Fisher information matrix (1) of GLM. An example of finding the Fisher information matrix for GLM is provided in Example S1.1 of Supplementary Material.
@@ 2.3  Fisher information for multinomial logistic model
We assume that the responses $\mathbf{Y}_{i}=(Y_{i1},\dots,Y_{iJ})$ follow a multinomial distribution with probabilities $(\pi_{i1},\dots,\pi_{iJ})$, where $\pi_{ij}=P(Y_{i}=j\mid{\mathbf{x}}_{i})$, $Y_{i}\in\{1,\ldots,J\}$ is the $i$th original categorical response, $j=1,\dots,J$, and $\sum_{j}\pi_{ij}=1$. If the response variables are nominal, in other words, there is no natural ordering among different levels, a commonly used MLM is the baseline-category logit model with the $J$th level selected as the baseline category. If the response variable is ordinal, that is, we have a natural ordering of response levels, there are three commonly used MLMs in the literature: cumulative logit model, adjacent logit model, and continuation-ratio logit model (Bu et al., 2020; Dousti Mousavi et al., 2023; Wang and Yang, ).
In addition to different logit models, the proportional odds (po) assumption is an important concept in MLMs. The po assumption is a parsimonious model assumption, where a model simultaneously uses all J1J-1 logits in a single model with the same coefficients. This means the covariate effect is constant on the odds ratio among different response levels. When the assumption doesn’t hold, the model is referred to as a non-proportional odds (npo) model and has more parameters in it. When the assumption only holds for part of the parameters, the model is referred to as a partial proportional odds (ppo) model. Commonly used multinomial logit models with po, npo, or ppo assumptions can be summarized in a unified matrix form (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020):
@@
$\mathbf{C}^{\top}=\begin{pmatrix}\mathbf{I}_{J-1}&-\mathbf{I}_{J-1}&\mathbf{0}_{J-1}\\ \mathbf{0}^{\top}_{J-1}&\mathbf{0}^{\top}_{J-1}&1\\ \end{pmatrix}$
@@ 3  Examples
Example Specific Finding $I$ set for exact allocation conversion in trial_data and trauma_data examples (see Section 3 and Section S4 in Supplementary Material). iset_func_trauma(), iset_func_trial()
@@ 3.1  Applications in paid research study: $\tt trial\_data$ example
The trial_data dataset is a simulated data for a toy example of paid research studies. This study includes a cohort of $500$ patients for a clinical trial, with gender and age as stratification factors. A logistic regression model incorporates these factors as covariates: gender (coded as 0 for female and $1$ for male) and age (coded as two dummy variables age_1age\_1 and age_2age\_2 with (age_1,age_2)=(0,0)(age\_1,age\_2)=(0,0) for age group $18\sim 25$; $(1,0)$ for age group $26\sim 64$, and $(0,1)$ for age group $65$ and above). For simplicity, the study assumes binary gender options and a tripartite age categorization. In total, there are $m=6$ combinations of covariate factors. In practice, non-binary gender options and a “prefer not to answer” choice may be included to respect gender diversity and protect patient confidentiality. The response $Y$ denotes the treatment’s efficacy (0 indicating ineffectiveness, $1$ indicating effectiveness). The data is generated by the logistic regression model
@@
Figure 4: The number of patients from different gender (F, M) and age groups (182518-25, 266426-64, and $\geq 65$) and their responses (0 indicating ineffectiveness, $1$ indicating effectiveness) to treatment in trial_data of CDsampling package.
@@
There are $m=6$ design points, corresponding to gender and age group combinations $(x_{gender\_i},$ $x_{age\_1i},x_{age\_2i})=(0,0,0)$, $(0,1,0)$, $(0,0,1)$, $(1,0,0)$, $(1,1,0)$, and $(1,0,1)$, respectively. The numbers of available volunteers in the six categories are $(N_{1},N_{2},\ldots,N_{6})=(50,40,10,$ $200,150,50)$. Our goal is to find the constrained D-optimal allocation $(w_{1},w_{2},\dots,w_{6})$ in the set of all feasible allocations $S=\{(w_{1},\ldots,$ $w_{m})^{T}\in S_{0}\mid nw_{i}\leq N_{i},i=1,\ldots,m\}$.
We consider the logistic regression model (3), to find the locally D-optimal design, we assume, for illustrative purposes, that the model parameters $(\beta_{0},\beta_{1},\beta_{21},\beta_{22})$ =(0,3,3,3)=(0,3,3,3). We may define the parameters and the model matrix as follows. Subsequently, we find the $\mathbf{W}$ matrix in (1) for the Fisher information matrix $\mathbf{F}$.
@@
  • $\mathbf{w}_{0}$: reports the random initial approximate allocation used to initialize optimization.
@@
In this example, the D-optimal exact allocation is to sample $50$ subjects from the “female, 182518-25” subgroup, $40$ subjects from the “female, 266426-64” subgroup, $10$ subjects from the “female, $\geq 65$” subgroup, and $100$ subjects from the “male, 182518-25” subgroup. Such a design may not explore all the design space, and lead to extreme design cases. In practice, allocating some subjects to the omitted subgroups “male, 266426-64” and “male, $\geq 65$” subgroups could improve the robustness and reduce the risk of overfitting.
Alternatively, one may aim for EW D-optimal allocations when partial coefficients information available with the $\mathbf{W}$} matrix substituted by the expectation (EW stands for expected weighted). To calculate these expectations, one may define prior distributions for the parameters based on available information. For instance, in this scenario, we assume the following independent prior distributions: $\beta_{0}\sim$ uniform(2,2)(-2,2), $\beta_{1}\sim$ uniform(1,5)(-1,5), $\beta_{21}\sim$ uniform(1,5)(-1,5), and $\beta_{22}\sim$ uniform(-1, 5). Subsequently, the diagonal elements of $\mathbf{W}$} are determined through integration. For $i=1,\dots,m$ and $\eta_{i}=\beta_{0}+x_{gender\_i}\beta_{1}+x_{age\_i1}\beta_{21}+x_{age\_i2}\beta_{22}$, we calculate the key component $E(\nu_{i})$ of the $i$th diagonal element of $\mathbf{W}$} through:
$\int_{-1}^{5}\int_{-1}^{5}\int_{-1}^{5}\int_{-2}^{2}\frac{\exp(\eta_{i})}{(1+\exp(\eta_{i}))^{2}}{\rm Pr}(\beta_{0}){\rm Pr}(\beta_{1}){\rm Pr}(\beta_{21}){\rm Pr}(\beta_{22})\,d\beta_{0}d\beta_{1}d\beta_{21}d\beta_{22}$
@@
> bounded_uniform(Ni=c(50, 40, 10, 200, 150, 50), nsample=200, label=label)Optimal Sampling Results:================================================================================Optimal exact allocation:           F, 18-25 F, 26-64 F, >=65 M, 18-25 M, 26-64 M, >=65allocation 38.0     38.0     10.0    38.0     38.0     38.0--------------------------------------------------------------------------------
@@
To perform a comparison analysis on different sampling strategies, including the constrained D-optimal allocation, the constrained EW D-optimal allocation with uniform priors, the constrained uniform allocation, simple random sample without replacement (SRSWOR), as well as the full data (all the $500$ patients enrolled), we simulate their responses $Y_{ij}$’s based on model (3) with parameter values $(0,3,3,3)$. We apply each sampling strategy to obtain a sample of $n=200$ observations out of $500$, and estimate the parameters using the $200$ observations. The exception is the full data method, where estimation is performed using all $500$ patients. We use the root mean square error (RMSE) to measure the accuracy of the estimates (see Section 4 of Huang et al. (2025) for more theoretical and technical details). We repeat the procedure $100$ times and display the corresponding RMSEs in Figure 5 and simulation codes in Supplementary Material Section S3 (Wickham, 2016; Wickham et al., 2023). Obviously, if we use the full data (all $500$ patients) to fit the model, its RMSE attains the lowest. Besides that, the constrained locally D-optimal allocation and the constrained EW D-optimal allocation have a little higher RMSEs than the full data estimates, but outperform SRSWOR and the constrained uniform allocation. The sampling strategy based on the constrained uniform allocation doesn’t need any model information and is a more robust sampling scheme, which is still better than SRSWOR.
@@ 3.2  Applications in paid research study: $\tt trauma\_data$ example
In this example, we have $m=8$ subgroups, which are combinations of the two covariates categories: dose levels and severity grades. We aim to enroll $n=600$ patients from the $802$ available patients. The collection of feasible allocations is inherently constrained by the number of patients in different severity grades, defined as $S=\{(w_{1},\ldots,w_{8})^{\top}\in S_{0}\mid n(w_{1}+w_{2}+w_{3}+w_{4})\leq 392,n(w_{5}+w_{6}+w_{7}+w_{8})\leq 410\}$. The constraints specify that in the sample, the number of patients with mild symptoms must not exceed $392$ across all dose levels, while those with moderate/severe symptoms must not exceed $410$.
@@
> J=5; p=12; m=8; #response levels; parameters; subgroups> #coefficients> beta = c(-4.047, -0.131, 4.214, -2.225, -0.376, 3.519, -0.302, -0.237,  2.420, 1.386,+ -0.120,  1.284)> #define design matrix of 8 subgroups> Xi=rep(0,J*p*m); dim(Xi)=c(J,p,m);> Xi[,,1] = rbind(c( 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+           c( 0, 0, 0, 1, 1, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 1, 1, 0, 0, 0, 0),+           c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,2] = rbind(c( 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+           c( 0, 0, 0, 1, 2, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 1, 2, 0, 0, 0, 0),+           c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,3] = rbind(c( 1, 3, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+           c( 0, 0, 0, 1, 3, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 1, 3, 0, 0, 0, 0),+           c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,4] = rbind(c( 1, 4, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+           c( 0, 0, 0, 1, 4, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 1, 4, 0, 0, 0, 0),+           c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,5] = rbind(c( 1, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+           c( 0, 0, 0, 1, 1, 1, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 1, 1, 1, 0, 0, 0),+           c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 1), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,6] = rbind(c( 1, 2, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+           c( 0, 0, 0, 1, 2, 1, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 1, 2, 1, 0, 0, 0),+           c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 1), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,7] = rbind(c( 1, 3, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+           c( 0, 0, 0, 1, 3, 1, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 1, 3, 1, 0, 0, 0),+           c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 1), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,8] = rbind(c( 1, 4, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+           c( 0, 0, 0, 1, 4, 1, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 1, 4, 1, 0, 0, 0),+           c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 1), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))
@@
As the trauma_data example doesn’t have bounded constraints, to find the constrained uniform sampling allocation, we use approxtoexact_constrained_func() with one subject in each stratum or subgroup as the input, that is, the approximate allocation $\mathbf{w}=(1/600,1/600,1/600,1/600,$ $1/600,1/600,1/600,1/600)$. The corresponding $I$ set function is provided in the CDsampling package, and it can be easily defined according to other constraints, see Section S4 of Supplementary Material. Note that the determinant provided by approxtoexact_constrained_func() for different designs are not comparable, as the criteria Fdet_func differ.